The expression x³-4x²+cx+d such that x + 1 is its factor, and its value is 1
when x is -2 . Find c and� d
1
Expert's answer
2019-05-22T05:09:18-0400

Since f(-1)=0 and f(-2)=1, make a system of equations:

f(-1)=-1-4-c+d=0

f(-2)=-8-16-2c+d=1

or

-c+d=5

-2c+d=25

Subtract the second equation from the first one: c=-20, and substitute this value for c into the first equation and get d=-15.

Answer: c=-20; d=-15